4.5 gm of a metallic chloride contains 35.5 gm of chlorine .then what is the equivalent mass of metal
 

Grade:12th pass

1 Answers

Charchit Tailong
8 years ago
For an element theequivalent weightis the quantity that combines with or replaces 1.00797 grams (g) of hydrogen or 7.9997 g of oxygen; or, theweightof an element that is liberated in an electrolysis (chemical reaction caused by an electric current) by the passage of 96,500 coulombs of electricity.
Equivalent weight (also known as gram equivalent) is a term which has been used in several contexts in chemistry. In its most general usage, it is the mass of one equivalent, that is the mass of a given substance which will:

* supply or react with one mole of hydrogen ions (H+) in an acid–base reaction; or
* supply or react with one mole of electrons in a redox reaction.[1]

Let us take ticl4(metal chloride)=4.5 gms
35.5gm chlorine

equivalent mass=46.5+35.5/4.5=82/4.5=18.22 gms answer